The length of a garden is 5 feet longer than its width. A 2 foot wide sidewalk
surrounds the garden. The area of the sidewalk is 76 sq. feet. What are the
dimensions of the garden?

Answer by mananth(16949)

The length of a garden is 5 feet longer than its width. A 2 foot wide sidewalk
surrounds the garden. The area of the sidewalk is 76 sq. feet. What are the
dimensions of the garden?
let the width of the garden be x
The length will be x+5 feet
The sidewalk width = 2 feet.
Width of garden + sidewalk = x+4 ( two on either side)
Length of garden + sidewalk = (x+5) +4
Total area = (x+9)(x+4)
Area of sidewalk = 76
(x+9)(x+4) - x(x+5) = 76
x^2+13x+36 - x^2-5x= 76
8x=76-36
8x = 40
x=5 feet width of the garden
5+5 = 10 feet is the length of the garden
Ananth